OpenStudy (anonymous):

Find the number of permutations of the first 8 letters of the alphabet taking 2 letters at a time.

OpenStudy (igreen):

\(\dfrac{n!}{(n - r)!}\) \(\dfrac{8!}{(8 - 3)!}\) Simplify, what's 8 - 3?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

336?

OpenStudy (igreen):

Yes, so we have: \(\dfrac{8!}{5!}\) Note, the exclamation points mean to multiply all the numbers below it. For example, 4! is equal to 4 * 3 * 2 * 1 Expand both 8! and 5!: \(\dfrac{8 \times 7 \times 6 \times 5 \times 4 \times 3 \times 2 \times 1}{5 \times 4 \times 3 \times 2 \times 1}\) \(\dfrac{8 \times 7 \times 6 \times \cancel{5 \times 4 \times 3 \times 2 \times 1}}{\cancel{5 \times 4 \times 3 \times 2 \times 1}}\) We're left with: \( 8 \times 7 \times 6\) Multiply: \(336\) You're correct!